There are three questions of 3D geometry,,can you plz solve these questions if you dont mind...


 


1.Find out the equation of the line passing through (1,1,1) and parallal to the plane 2x+3y+z+5=0


 


2.If the plane (x/2)+(y/3)+(z/4)=1 cuts the coordinates axes in A,B,C then the area of Δ ABC is _____________


 


3.The equation of a plane through the line of intersection of 2x+3y+z-1=0 and x+5y-2z+7=0 and parallel to the line y=0=z is _______________

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

 let the equation of the line passing through (1,1,1) and parallal to the plane 2x+3y+z+5=0  is

(x-1)/a  = (y-1)/b = (z-1)/c

wher a,b,c are the direction ratios of line.

and given direction ratios  of the normal of the plane is 2,3,1

so a*2 + b*3 + c*1 =0

for different combination of a,b,c different line are posible so infinite lines are possible.

for example for c=1 a=1 b=-1 one line is possible

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin